RC Meeting and Run Plan, Friday, February 22
RC Meeting Information
RC Daily Meetings are at 4:30 PM, 2nd floor Counting House Meeting Room
- Bluejeans remote connection information:
To join the Meeting: https://bluejeans.com/861439920 To join via Room System: Video Conferencing System: bjn.vc -or-126.96.36.199 Meeting ID : 861439920 To join via phone : 1) Dial: +1.408.740.7256 (US (San Jose)) +1.888.240.2560 (US Toll Free)) +1.408.317.9253 (US (Primary, San Jose)) (see all numbers - http://bluejeans.com/numbers) 2) Enter Conference ID : 861439920
- Continued running J/Psi on kinematic setting 3 high-t
- SHMS HB problems right after last meeting. Called Steve. Jack came to help remote reboot of the PLC. This seemed to fix things.
- Strange target alarms. Loss of cooling power?
- Running at 55 -60 uA. Will ask for more if beam is stable.
- Replace camera 3 (temporarily). Needs better cable placement.
- Might turn off camera rotation. Need to follow up with Steve.
Current Run Plan
- J/Psi Experiment:
- Production Phase 3: high-t
Notices in Effect for Shift Crew
- Shift Leader is responsible for 50k replays. Even if the TO is actually doing them.
- Target Operators: When the alarm handler goes white, it is disconnected from the IOC. Log it and include email@example.com every time. Acknowledge the warning. If it is still white and not coming back, restart the alarm handler. Do not permanently silence the alarm handler.
- Missing Reference Times during the replay: check in the 50k. Check in full replay, if able. If it happens, call RC. Guidance to solve it is here: https://logbooks.jlab.org/entry/3654584
- Target camera is inoperable at the moment.
- Need to notify Steve to turn on rotation controller before rotation.